### v3

# The $C^{0,1}$ Itô-Ventzell formula for weak Dirichlet processes

Abstract
This paper proves an extension of the Itô-Ventzell formula that applies to stochastic flows in $C^{0,1}$ for continuous weak Dirichlet processes. We apply this theorem, for example, to give a representation result for strong solutions of time-dependent elliptic SPDEs, to derive formulas for quadratic variations, and to relax assumptions in a financial mathematics context.
Keywords: Generalized Itô-Ventzell formula, weak Dirichlet process, Integration via regularization
2020 MSC: 60H05, 60G48, 60G44, 60J65
1 Introduction
In this paper, we show an extension of the Itô-Ventzell formula for continuous weak Dirichlet processes, such that the underlying stochastic flow only needs to be in $C^{1}$ instead of being in $C^{2}$ in the spacial argument under an additional mild assumption. The Itô-Ventzell formula is a generalization of the Itô formula for stochastic flows. Specifically, let $X=M+A$ be a continuous semimartingale, where $M$ is a local martingale and $A$ is a process of finite variation, and let $F(t,x)$ be a stochastic flow with local characteristics $(\beta(t,x),\gamma(t,x))$, i.e., $F(t,x)$ is a random field and admits the Itô dynamics
$F(t,x)=F(0,x)+\int_{0}^{t}\beta(r,x)\mathrm{d}r+\int_{0}^{t}\gamma(r,x)\mathrm{d}W_{r},$
where $W$ is a Brownian Motion. Moreover, we assume that $F$ is a continuous progressively measurable process for fixed $x$ and twice differentiable in $x$ for fixed $t$ and $\omega$. Note that the dependence of $F$, $\beta$, and $\gamma$ on $\omega$ is suppressed. Then the Itô-Ventzell formula states that $F(t,X_{t})$ is also a continuous semimartingale and has the following decomposition:
$F(t,X_{t})=$ $\ F(0,X_{0})+\int_{0}^{t}\gamma(s,X_{s})\mathrm{d}W_{s}+\int_{0}^{t}F_{x}(s,X_{s})\mathrm{d}M_{s}+\mathcal{B}^{X}(F)_{t},$
where
$\mathcal{B}^{X}(F)_{t}=\int_{0}^{t}\beta(s,X_{s})\mathrm{d}s+\int_{0}^{t}F_{x}(s,X_{s})\mathrm{d}A_{s}+\displaystyle\frac{1}{2}\int_{0}^{t}F_{xx}(s,X_{s})\mathrm{d}[X]_{s}+\int_{0}^{t}\gamma_{x}(s,X_{s})\mathrm{d}[X,W]_{s}.$
The Itô-Ventzell formula links stochastic processes to stochastic partial differential equations (SPDEs) and has various applications. It is used, for instance, to analyze various optimal control problems, to determine localization errors of SPDEs, for generalized stochastic forward integrals, for forward performance processes, and to develop pricing models. We apply our new version of the Itô-Ventzell formula to derive stochastic calculus formulas and to obtain a representation result for strong solutions of time-dependent stochastic elliptic PDEs. Here, we generalize naturally the solution theory of SPDEs to the setting of the paper. Moreover, we apply our result in a financial mathematics context to obtain the dynamics of the portfolio of a large investor whose trading influences prices.
Multiple works generalize the classical Itô formula in different directions, see, for instance, Föllmer et al. [12], Elworthy et al. [9] or da Prato et al. [7]. We focus on the extension discovered by Ventzell [26] in 1965, the Itô-Ventzell formula, which was generalized later by Kunita [17]. Since then, multiple works extended the Itô-Ventzell formula, see, for instance, Flandoli and Russo [11], Krylov [16], Karachanskaya [14], Kohlmann et al. [15], or dos Reis and Platonov [8].
Our results hold for weak Dirichlet processes, the sum of a local martingale and an adapted process with zero covariation with all continuous local martingales, including semimartingales as a special case. Since the classical theorem of stochastic integration is defined only with respect to semimartingales, we use an extension of stochastic integration developed by Russo and Vallois [20, 21, 22, 22, 23, 24], Coquet et al. [6], and Bandini and Russo [1, 2]. Considering weak Dirichlet processes is not unusual in generalizations of Itô’s-formula which require milder smoothness conditions on the function $F$, see Gozzi and Russo [13] who generalized the classical Itô formula to $C^{0,1}$, Bouchard et al. [4] for the functional Itô formula, or Bouchard and Vallet [5] for the Itô-Dupire formula.
The paper is structured as follows: In Section 2, we introduce weak Dirichlet processes, generalized stochastic integration, and stochastic flows. Section 3 gives our result, namely the extension of the Itô-Ventzell formula to stochastic flows in $C^{0,1}$, while the proof of this extension is shown in Section 4. Section 5 gives some applications of our extended version of the Itô-Ventzell formula.
2 Basic Definitions
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ be a stochastic basis where the filtration $\mathcal{F}$ is generated by a Brownian Motion $W$ satisfying the usual conditions and $[0,T]$ is the observation period. We define $f_{x}(t,x)$ to be the partial derivative of $f$ in the second argument and denote higher derivatives similarly. In this paper, we use the space of progressively measurable and continuous processes equipped with the topology of uniform convergence on compacts in probability. We denote by $C^{n,m+\alpha}$ the set of functions which are $n$-times continuously differentiable in the first and $m$-times in the second argument where the $m$th derivative is $\mathrm{d}\mathbb{P}\times\mathrm{d}t$-a.s. locally $\alpha$-Hölder continuous with $\alpha>0$. If the $m$th derivative is just continuous, we simply write $C^{n,m}$, and $C^{0,0}$ denotes the set of all continuous functions. Furthermore, we denote by $\mathcal{L}^{1}(\mu)$ (resp. $\mathcal{L}^{2}(\mu)$) the space of integrable (resp. square integrable) random variables/measurable functions with respect to a measure $\mu$.
Definition 2.1.
The sequence of progressively measurable processes $(H^{n})$ converges uniformly on compacts in probability ($ucp$) to a process $H$ if $\sup_{s\in[0,t]}\left|H^{n}_{s}-H_{s}\right|\xrightarrow{n\to\infty}0$ in probability for all $t>0$.
Our processes may not be semimartingales and may not admit a quadratic (co-)variation. For this purpose, we need generalizations of these concepts, see Gozzi and Russo [13, p.1567]:
Definition 2.2.
Let $X$ and $Y$ be two continuous processes of the same dimension. Then, we define for all $t\in[0,T]$:
$[X,Y]_{t}$ $=\lim\limits_{\varepsilon\downarrow 0}\int_{0}^{t}\displaystyle\frac{(X_{r+\varepsilon}-X_{r})(Y_{r+\varepsilon}-Y_{r})}{\varepsilon}\mathrm{d}r,$
$\int_{0}^{t}X_{r}\mathrm{d}^{-}Y_{r}$ $=\lim\limits_{\varepsilon\downarrow 0}\int_{0}^{t}X_{r}\displaystyle\frac{Y_{r+\varepsilon}-Y_{r}}{\varepsilon}\mathrm{d}r,$
if those limits exist in the sense of $ucp$-convergence. We call $[X,Y]_{t}$ the covariation, and $\int_{0}^{t}X_{r}\mathrm{d}^{-}Y_{r}$ the forward integral. We use the short notation for the covariation: $[X]_{t}:=[X,X]_{t}$.
Definition 2.2 ensures that the covariation and the forward integral are again continuous processes and that the covariation is bilinear and symmetric. In particular, the (forward) integral against the covariation is well-defined. For a more detailed explanation of the covariation and the integral against it, see Russo and Vallois [22, pp. 83-97].
Remark 2.3.
Due to Definition 2.2, the functions inside of the covariation and the integral need to be defined on $[0,T+\varepsilon]$ instead of $[0,T]$. Consequently, we use the convention that these processes or functions are constantly continued on $(T,T+\varepsilon]$, e.g., we define $X_{r}=X_{T}$ for all $r>T$.
Definition 2.2 ensures that the covariation and the forward integral are continuous. If $X$ and $Y$ are semimartingales, then the concepts above coincide with the quadratic covariation and the usual notion of a stochastic integral. While semimartingales are of finite quadratic variation, this is, of course, not true for general continuous processes. Hence, we define the following:
Definition 2.4.
Let $X$ be a continuous and adapted process. If $[X]_{t}<\infty$ for all $t\in[0,T]$, then we call $X$ a finite quadratic variation (FQV) process.
The following definition is standard, see, for instance, Russo and Vallois [20, 21, 22, 22, 23, 24], Coquet et al. [6], Bandini and Russo [1, 2] or Gozzi and Russo [13, p.1570]:
Definition 2.5.
A real continuous and adapted process $X$ is called a weak Dirichlet process if $X$ is decomposable into a local martingale $M$ and a weak zero energy process $A$, i.e., $X=M+A$.
For the process $A$, it holds: $A_{0}=0$ and $[A,N]_{t}=0$ for all continuous local martingales $N$ and $t\in[0,T]$.
In general, this decomposition is not unique. However, when $M$ is required to be continuous, there exists a unique decomposition. In the following, we always use the latter decomposition. In particular, $A$ is continuous, since by assumption $X$ and $M$ are. If $X$ is additionally a semimartingale, $A$ is a continuous finite variation process.
Example 2.6.
The following processes are weak-Dirichlet processes according to Gozzi and Russo [13, pp.1570-1571], and Errami and Russo [10, p.267]:
  • all semimartingales,
  • all Dirichlet processes $X$, i.e., $X$ is decomposable into a local martingale and a zero quadratic variation process,
  • the process $X_{t}=\int_{0}^{t}G(t,r)\mathrm{d}M_{r}$, where $G:\{0\leq r\leq t\leq 1\}\to\mathbb{R}$ is a continuous $(\mathcal{F}_{r})$-measurable random field and $M$ is a local continuous $(\mathcal{F}_{t})$-martingale.
Contrary to the classical Itô formula, the Itô-Ventzell formula considers stochastic flows, which are defined in the following definition. Similar, but slightly different definitions can be found, e.g., in Kohlmann et al. [15, p.7] or dos Reis and Platonov [8, p.6].
Definition 2.7.
Let $F(t,x)=F(\omega,t,x):\Omega\times[0,\infty)\times\mathbb{R}^{d}\to\mathbb{R}$ be a stochastic flow with local characteristics $(\beta(t,x),\gamma(t,x))$, i.e., $F$ is a random field which admits the Itô dynamics
$F(t,x)=F(0,x)+\int_{0}^{t}\beta(r,x)\mathrm{d}r+\int_{0}^{t}\gamma(r,x)\mathrm{d}W_{r},$ (2.1)
where
  1. (a)
    $W$ is the Brownian motion which generates the filtration $\mathcal{F}$,
  2. (b)
    $F(\cdot,x)$ is a progressively measurable process for all fixed $x\in\mathbb{R}^{d}$,
  3. (c)
    $\beta:\Omega\times\mathbb{R}_{\geq 0}\times\mathbb{R}^{d}\to\mathbb{R}$ and $\gamma:\Omega\times\mathbb{R}_{\geq 0}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ are previsible random fields,
  4. (d)
    $\gamma(t,x)$ is $\mathrm{d}\mathbb{P}\times\mathrm{d}t$-a.s. locally $\alpha$-Hölder continuous in $x$ for an $\alpha>0$, and
  5. (e)
    for any $K\subset\mathbb{R}^{d}$ compact and for any $S\geq 0$ it holds a.s. that $\int_{0}^{S}\sup_{x\in K}|\beta(t,x)|\mathrm{d}t<\infty$, and $\int_{0}^{S}\sup_{x\in K}|\gamma(t,x)|^{2}\mathrm{d}t<\infty$.
The dependence on $\omega$ will be suppressed throughout this paper unless it is necessary for the steps of a proof.
Remark 2.8.
The local $\alpha$-Hölder continuity of $\gamma$ is necessary to ensure through Kolmogorov’s continuity theorem that the stochastic integral $\int_{0}^{t}\gamma(r,x)\mathrm{d}W_{r}$ is a.s. well-defined for all $(t,x)$. We refer to Kunita [17] for details.
3 Main Result
The following result states the Itô-Ventzell formula for stochastic flows in $C^{0,1}$:
Theorem 3.1.
Let $X$ be a continuous weak Dirichlet FQV process and $F\in C^{0,1}$ be a stochastic flow with local characteristics $(\beta(t,x),\gamma(t,x))$. Furthermore, we assume:
  1. 1a
    $\int_{0}^{T}\sup_{s\in[0,T]}|\beta(s,X_{t})|\mathrm{d}t<\infty$ a.s., or
  2. 1b
    there exist a $B_{1}>0$ such that $\int_{0}^{T}\sup_{s\in[0,T]}\mathbb{E}\left[\left|\beta(s,X_{t})\right|\right]\mathrm{d}t\leq B_{1}$.
  3. 2
    for all $K\subset\mathbb{R}^{d}$ compact and $p\geq 1$, we suppose the existence of a $C_{p}\in\mathbb{R}$ such that for any $x,y\in K$ holds that $\mathbb{E}\left[\sup_{t\in[0,T]}|\gamma(t,x)-\gamma(t,y)|^{p}\right]\leq C_{p}|x-y|^{p}$ and $\mathbb{E}\left[\sup_{t\in[0,T]}|\gamma(t,0)|^{p}\right]<\infty$.
  4. 3
    $\gamma(t,X_{t})$ is a.s. càdlàg (right continuous with left limits).
Then it holds for all $t\in[0,T]$:
$F(t,X_{t})=F(0,X_{0})+\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}M_{r}+\mathcal{B}^{X}(F)_{t},$
where $\mathcal{B}^{X}$ is linear and continuous (in $F$) with respect to the topology of $ucp$ convergence and $F(t,X_{t})$ is again a continuous weak Dirichlet process. Moreover,
  1. (a)
    $\mathcal{B}^{X}$ is a weak zero energy process, and
  2. (b)
    if $F(t,x)\in C^{0,2}$, $\beta(t,x)\in C^{0,1}$, and $\gamma(t,x)\in C^{0,1+\alpha}$ for an $\alpha\in(0,1)$, then it holds:
    $\mathcal{B}^{X}(F)_{t}=$ $\int_{0}^{t}\beta(r,X_{r})\mathrm{d}r+\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\gamma_{x}(r,X_{r})\mathrm{d}[X,W]_{r}+\frac{1}{2}\int_{0}^{t}F_{xx}(r,X_{r})\mathrm{d}[X]_{r}.$
In the last two integrals of Theorem 3.1 (b) both the integrand and the integrator are matrices. These integrals are, as usual, defined as the trace of the matrix of the corresponding integrals. The assumptions of Theorem 3.1 are satisfied (through Assumption 1a) if the mappings $(t,x)\mapsto\beta(t,x)$ and $(t,x)\mapsto\gamma(t,x)$ are a.s. continuous, $\sup_{t\in[0,T]}|\gamma(t,0)|$ is $p$-integrable for all $p\in[1,\infty)$ and $x\mapsto\gamma(\cdot,x)$ is a.s. locally Lipschitz continuous with a Lipschitz constant independent of $t$ which is integrable in $\omega$.
The theorem generalizes the result of Gozzi and Russo [13, pp.1571-1574] to stochastic flows.
Corollary 3.2.
Let the assumptions of Theorem 3.1 hold. Moreover, assume that $F(\cdot,X)$ is a local martingale. Then, it holds for all $t\in[0,T]$:
$F(t,X_{t})=F(0,X_{0})+\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}M_{r}.$
Proof. Using Theorem 3.1, we get:
$\mathcal{B}^{X}(F)_{t}=F(t,X_{t})-F(0,X_{0})-\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}-\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}M_{r}.$
Since $F(\cdot,X)$, $M$, and $W$ are local martingales, $\mathcal{B}^{X}(F)$ is a local martingale as well, and Theorem 3.1 states that $\mathcal{B}^{X}(F)$ is a continuous weak zero energy process. Thus, $[\mathcal{B}^{X}(F)]=0$ which implies that $\mathcal{B}^{X}(F)\equiv 0$. Hence, Theorem 3.1 finalizes the proof.
Remark 3.3.
Theorem 3.1 in Krylov [16] provides a distribution-valued version of the Itô-Ventzell formula. The main difference between his and our version of the Itô-Ventzell formula is that in Theorem 3.1 in Krylov [16], one needs the existence of a second derivative in the space direction (in the distributional sense) of the function $F$, while we only need the first derivative. On the other hand, Krylov [16] considers possibly infinitely many space directions and does not need joint continuity in time and space directions and the previsibility of the characteristics. Contrary, we need less regularity for the characteristics, i.e., we do not need to assume that $\beta$ and $\gamma$ are continuous. Regarding the integrability, our Theorem 3.1 and Theorem 3.1 in Krylov [16] have somewhat similar assumptions.
4 Proof of Theorem 3.1
Before proving the theorem, we need some preliminary results. First of all, we assume $d=1$ to simplify the notation. In the whole section, we use the convention from Remark 2.3 for the constant continuation of processes and functions.
By recalling Proposition 1.2 of Russo and Vallois [22, p.89] it holds:
Lemma 4.1.
Let $X$ and $Y$ be two FQV processes and $H$ be a càdlàg process defined on the interval $[0,T]$. Then, it holds:
$\int_{0}^{t}H(r)\frac{(X_{r+\varepsilon}-X_{r})(Y_{r+\varepsilon}-Y_{r})}{\varepsilon}\mathrm{d}r\xrightarrow[ucp]{\varepsilon\to 0}\int_{0}^{t}H(r)\mathrm{d}[X,Y]_{r}$
Lemma 4.2.
Let $X$, $Y$ be two FQV processes, $t\in[0,T]$. Moreover, let $Z^{\varepsilon}$ be a measurable process in $r$ for all $\varepsilon>0$ such that $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$. Then:
$\int_{0}^{t}Z^{\varepsilon}_{r}\frac{(X_{r+\varepsilon}-X_{r})(Y_{r+\varepsilon}-Y_{r})}{\varepsilon}\mathrm{d}r\xrightarrow[ucp]{\varepsilon\to 0}0.$
Proof. First, we note that $Z^{\varepsilon}$ is a.s. bounded for all $\varepsilon>0$ (with a bound possibly depending on $\omega$), i.e., the integral is well-defined. Then, we get:
$\sup_{t\in[0,T]}\left|\int_{0}^{t}Z^{\varepsilon}_{r}\frac{(X_{r+\varepsilon}-X_{r})(Y_{r+\varepsilon}-Y_{r})}{\varepsilon}\mathrm{d}r\right|$ $\leq\int_{0}^{T}\left|Z^{\varepsilon}_{r}\frac{(X_{r+\varepsilon}-X_{r})(Y_{r+\varepsilon}-Y_{r})}{\varepsilon}\right|\mathrm{d}r$
$\leq\sup_{r\in[0,T]}\left|Z^{\varepsilon}_{r}\right|\cdot\int_{0}^{T}\left|\frac{(X_{r+\varepsilon}-X_{r})(Y_{r+\varepsilon}-Y_{r})}{\varepsilon}\right|\mathrm{d}r,$
and the lemma follows from Lemma 4.1.
For the following lemma, we need the concept of boundedness in probability, see, e.g., van der Vaart [25]:
Definition 4.3.
Let $I$ be an index set. A set of random variables $(X_{\delta})_{\delta\in I}$ is bounded in probability (also known as uniformly tight) if there exists for all $\varepsilon>0$ an $M>0$ such that $\sup_{\delta\in I}\mathbb{P}(|X_{\delta}|>M)<\varepsilon$.
Note that $\mathcal{L}^{1}$ boundedness implies via Markov’s inequality boundedness in probability.
Lemma 4.4.
Let $X$ be a continuous FQV process, $t\in[0,T]$, and $Z^{\varepsilon}$ be a measurable and a.s. bounded process for all $\varepsilon>0$ such that $\int_{0}^{T}|Z_{r}^{\varepsilon}|\mathrm{d}r$ is bounded in probability in $\varepsilon>0$. Then:
$\int_{0}^{t}Z^{\varepsilon}_{r}\left(X_{r+\varepsilon}-X_{r}\right)\mathrm{d}r\xrightarrow[ucp]{\varepsilon\to 0}0.$
Proof. Let $\omega\in\Omega$ be arbitrary such that $X(\omega)$ is continuous. Thus, $X(\omega)$ is uniformly continuous on $[0,T]$, i.e., $\sup_{r\in[0,T]}\left|X_{r+\varepsilon}(\omega)-X_{r}(\omega)\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$. Hence, we get:
$\sup_{t\in[0,T]}\left|\int_{0}^{t}Z^{\varepsilon}_{r}\left(X_{r+\varepsilon}-X_{r}\right)\mathrm{d}r\right|$ $\leq\int_{0}^{T}\left|Z^{\varepsilon}_{r}\left(X_{r+\varepsilon}-X_{r}\right)\right|\mathrm{d}r$
$\leq\int_{0}^{T}\left|Z^{\varepsilon}_{r}\right|\mathrm{d}r\cdot\sup_{r\in[0,T]}\left|X_{r+\varepsilon}(\omega)-X_{r}(\omega)\right|\xrightarrow[\mathbb{P}]{\varepsilon\to 0}0.$
Lemma 4.5.
Let $X$ be a FQV process, $t\in[0,T]$, and $Z^{\varepsilon}$ be a measurable and a.s. bounded process for all $\varepsilon>0$ such that $\int_{0}^{T}\left|Z_{r}^{\varepsilon}\right|^{2}\mathrm{d}r\xrightarrow[\mathbb{P}]{\varepsilon\to 0}0$. Then:
$\int_{0}^{t}Z^{\varepsilon}_{r}\frac{X_{r+\varepsilon}-X_{r}}{\sqrt{\varepsilon}}\mathrm{d}r\xrightarrow[ucp]{\varepsilon\to 0}0.$
Proof. We get with Cauchy-Schwarz:
$\sup_{t\in[0,T]}\left|\int_{0}^{t}Z^{\varepsilon}_{r}\frac{X_{r+\varepsilon}-X_{r}}{\sqrt{\varepsilon}}\mathrm{d}r\right|\leq\int_{0}^{T}\left|Z^{\varepsilon}_{r}\frac{X_{r+\varepsilon}-X_{r}}{\sqrt{\varepsilon}}\right|\mathrm{d}r\leq\sqrt{\int_{0}^{T}\left|Z_{r}^{\varepsilon}\right|^{2}\mathrm{d}r}\sqrt{\int_{0}^{T}\frac{(X_{r+\varepsilon}-X_{r})^{2}}{\varepsilon}\mathrm{d}r}.$
The first factor converges in probability to 00 for $\varepsilon\to 0$ due to the assumption of the lemma. The second factor converges in probability to $\sqrt{[X]_{T}}$ for $\varepsilon\to 0$ by Definition 2.2.
Lemma 4.6.
In the setting of Theorem 3.1, it holds for all $t\in[r,T]$ and for all $\mathcal{F}_{r}$-measurable random variables $G$ such that $\int_{0}^{t}|\beta(s,G)|\mathrm{d}s<\infty$ a.s. and $\int_{0}^{t}|\gamma(s,G)|^{2}\mathrm{d}s<\infty$ a.s.:
$\left.\int_{r}^{t}\beta(s,x)\mathrm{d}s\right|_{x=G}$ $=\int_{r}^{t}\beta(s,G)\mathrm{d}s,$ $\left.\int_{r}^{t}\gamma(s,x)\mathrm{d}W_{s}\right|_{x=G}$ $=\int_{r}^{t}\gamma(s,G)\mathrm{d}W_{s}.$
Proof. The first property follows directly from the $\omega$-wise definition of the “$\mathrm{d}s$”-integral. The second property is an immediate consequence of Lemma 3.3 in Russo and Vallois [23].
Lemma 4.7.
Consider the setting of Theorem 3.1 and assume that $X$ takes only values in a compact set $K\subset\mathbb{R}$. Then Assumption 2 implies that there exist $B_{2}>0$ such that $\sup_{s\in[0,T]}\mathbb{E}\left[\gamma(s,X_{t})^{2+\varrho}\right]\leq B_{2}$ for all $t\in[0,T]$ and for all $\varrho>0$.
Proof. The lemma follows from Assumption 2 in Theorem 3.1.
Now, we are ready to prove our main result:
  • Proof of Theorem 3.1.
    We can write $\mathcal{B}^{X}$ as
    $\mathcal{B}^{X}(F)_{t}=F(t,X_{t})-F(0,X_{0})-\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}-\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}M_{r}.$
    Hence, $\mathcal{B}^{X}$ is linear and continuous (in $F$) with respect to the $ucp$ topology. The property that $F(t,X_{t})$ is a weak Dirichlet process follows directly from $\mathcal{B}^{X}$ being a weak zero energy process (which we show later) since $M$ and $W$ are local martingales.
    If $F(t,x)\in C^{0,2}$ and $\gamma(t,x),\beta(t,x)\in C^{0,1}$ with $\alpha$-Hölder-continuous derivatives in $x$, then Flandoli and Russo [11, p.279] give us the Itô-Ventzell formula for weak Dirichlet processes:
    $F(t,X_{t})=$ $F(0,X_{0})+\int_{0}^{t}\beta(r,X_{r})\mathrm{d}r+\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}X_{r}$
    $+\int_{0}^{t}\gamma_{x}(r,X_{r})\mathrm{d}[X,W]_{r}+\frac{1}{2}\int_{0}^{t}F_{xx}(r,X_{r})\mathrm{d}[X]_{r}.$
    Next note that $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}X_{r}$ exists. Due to $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}M_{r}=\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}M_{r}$ coinciding with the classical stochastic integral, $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}$ also exists. Moreover, when carefully scanning the to our setting adapted proof of the Itô-Ventzell formula in Flandoli and Russo [11, pp.279-280], we can reduce the regularity assumptions of $\beta$ and $\gamma$ due to Lemma 4.6. Thus, it is sufficient to assume that $\beta(t,x)\in C^{0,1}$ (without Hölder-continuity of $\beta_{x}(t,x)$ in $x$) and $\gamma(t,x)\in C^{0,1+\alpha}$ (with only local $\alpha$-Hölder-continuity of $\gamma_{x}(t,x)$ in $x$) in case (b).
    Hence, it remains to prove the case (a) where $F\in C^{0,1}$ and the Assumptions 1-3 hold:
    First of all, since $X$ is continuous, we can use a classical localization argument and assume that there exists a compact set $K\subset\mathbb{R}$ such that $X_{t}(\omega)\in K$ for all $t\in[0,T]$ and almost all $\omega\in\Omega$. Then, by using again classical localization arguments, we can assume that $\sup_{x\in K}|\beta(t,x)|\in\mathcal{L}^{1}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$, $\sup_{x\in K}|\gamma(t,x)|\in\mathcal{L}^{2}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$ both from Definition 2.7(e). Now, the prerequisites of Lemma 4.6 are fulfilled, and thus all of the (stochastic) integrals in this proof are well-defined.
    Let $N$ be a continuous local martingale. Since $N$, $M$, and $W$ are semimartingales, we know:
    $\left[\int_{0}^{\cdot}F_{x}(r,X_{r})\mathrm{d}M_{r},N\right]$ $=\int_{0}^{\cdot}F_{x}(r,X_{r})\mathrm{d}\left[M,N\right]_{r},$
    $\left[\int_{0}^{\cdot}\gamma(r,X_{r})\mathrm{d}W_{r},N\right]$ $=\int_{0}^{\cdot}\gamma(r,X_{r})\mathrm{d}\left[W,N\right]_{r}.$
    Hence, it remains to show:
    $\left[F(\cdot,X),N\right]=\int_{0}^{\cdot}F_{x}(r,X_{r})\mathrm{d}\left[M,N\right]_{r}+\int_{0}^{\cdot}\gamma(r,X_{r})\mathrm{d}\left[W,N\right]_{r}.$
    Using Definition 2.2, we consider the $ucp$ limit of
    $\int_{0}^{t}$ $\left[F(r+\varepsilon,X_{r+\varepsilon})-F(r,X_{r})\right]\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r$
    $=\int_{0}^{t}\left[F(r+\varepsilon,X_{r+\varepsilon})-F(r+\varepsilon,X_{r})\right]\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r+\int_{0}^{t}\left[F(r+\varepsilon,X_{r})-F(r,X_{r})\right]\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r$
    $=:I_{1}(t,\varepsilon)+I_{2}(t,\varepsilon).$
    We first show: $\lim_{\varepsilon\to 0}I_{1}(t,\varepsilon)=\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}\left[M,N\right]_{r}$. It is
    $I_{1}(t,\varepsilon)$ $=\int_{0}^{t}F_{x}(r,X_{r})\left(X_{r+\varepsilon}-X_{r}\right)\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r$
    $\hskip 12.0pt+\int_{0}^{t}\left(F_{x}(r+\varepsilon,X_{r})-F_{x}(r,X_{r})\right)\left(X_{r+\varepsilon}-X_{r}\right)\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r$
    $\hskip 12.0pt+\int_{0}^{t}\int_{0}^{1}\left\{F_{x}(r+\varepsilon,X_{r}+\lambda(X_{r+\varepsilon}-X_{r}))-F_{x}(r+\varepsilon,X_{r})\right\}\mathrm{d}\lambda\left(X_{r+\varepsilon}-X_{r}\right)\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r$
    $=:I_{1,1}(t,\varepsilon)+R_{1,1}(t,\varepsilon)+R_{1,2}(t,\varepsilon).$
    We know that $r\mapsto F_{x}(r,X_{r})$ is càdlàg since $F\in C^{0,1}$ and $X$ is a.s. continuous. Thus, Lemma 4.1 implies that $\lim_{\varepsilon\to 0}I_{1,1}(t,\varepsilon)=\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}\left[X,N\right]_{r}=\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}\left[M,N\right]_{r}$ since $[A,N]_{r}\equiv 0$. Hence, it remains to show: $\lim_{\varepsilon\to 0}R_{1,1}(t,\varepsilon)=0$ and $\lim_{\varepsilon\to 0}R_{1,2}(t,\varepsilon)=0$ in $ucp$. For this purpose, we define for $r\in[0,T]$ and $\varepsilon>0$:
    $Y_{r}^{\varepsilon}$ $:=F_{x}(r+\varepsilon,X_{r})-F_{x}(r,X_{r}),$
    $Z_{r}^{\varepsilon}$ $:=\int_{0}^{1}\left\{F_{x}(r+\varepsilon,X_{r}+\lambda(X_{r+\varepsilon}-X_{r}))-F_{x}(r+\varepsilon,X_{r})\right\}\mathrm{d}\lambda.$
    Let $\omega\in\Omega$ be arbitrary such that $r\to X_{r}(\omega)$ is continuous. Hence, there exists a convex compact set $K$ such that $X_{r}(\omega)\in K$ for all $r\in[0,T]$. Since $F_{x}(t,x)$ is continuous, we know that $F_{x}(t,x)$ is uniformly continuous on $[0,T]\times K$. Thus, for all $L>0$ there exists a $0<\rho_{\omega}\leq 1$ such that it holds for all $r\in[0,T]$ and for all $0<\varepsilon\leq\rho_{\omega}$: $\left|Y_{r}^{\varepsilon}(\omega)\right|\leq L$. (Note that $\left\lVert(r+\varepsilon,X_{r}(\omega))-(r,X_{r}(\omega))\right\rVert_{2}=\varepsilon\leq\rho_{\omega}$.) In particular, we get for any arbitrary $L>0$ a $0<\rho_{\omega}\leq 1$ such that $\sup_{r\in[0,T]}\left|Y_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$. Hence, $\sup_{r\in[0,T]}\left|Y_{r}^{\varepsilon}(\omega)\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$.
    Next, note that it holds that $X_{r}(\omega)+\lambda(X_{r+\varepsilon}(\omega)-X_{r}(\omega))\in K$ for all $\lambda\in[0,1]$ since $X_{r}(\omega),\linebreak X_{r+\varepsilon}(\omega)\in K$. Moreover, $X_{r}(\omega)$ is uniformly continuous in $r$ for $r\in[0,T]$, i.e., for all $\tilde{L}>0$ there exists a $0<\tilde{\rho}_{\omega}\leq 1$ such that for all $r\in[0,T]$ and for all $0<\varepsilon\leq\tilde{\rho}_{\omega}$ it holds: $\left|X_{r+\varepsilon}(\omega)-X_{r}(\omega)\right|\leq\tilde{L}$. Let $L>0$ arbitrary. Since $F_{x}(t,x)$ is uniformly continuous on $[0,T]\times K$, there exists a $\delta>0$ such that for all $x,y\in K$ with $\left\lVert(t,x)-(t,y)\right\rVert_{2}\leq\delta$ it holds that $\left|F_{x}(t,x)-F_{x}(t,y)\right|\leq L$ for all $t\in[0,T]$ fixed. Now choose $\tilde{L}=\delta$ and define $\rho_{\omega}:=\min\{\delta,\tilde{\rho}_{\omega}\}$. Then
    $\left\lVert(r+\varepsilon,X_{r}(\omega)+\lambda(X_{r+\varepsilon}(\omega)-X_{r}(\omega)))-(r+\varepsilon,X_{r}(\omega))\right\rVert_{2}=\lambda|X_{r+\varepsilon}(\omega)-X_{r}(\omega)|\leq\delta$
    for all $0<\varepsilon\leq\rho_{\omega}$, all $\lambda\in[0,1]$, and all $r\in[0,T]$. Hence, it holds that
    $\left|F_{x}(r+\varepsilon,X_{r}(\omega)+\lambda(X_{r+\varepsilon}(\omega)-X_{r}(\omega)))-F_{x}(r+\varepsilon,X_{r}(\omega))\right|\leq L$
    for all $0<\varepsilon\leq\rho_{\omega}$, all $\lambda\in[0,1]$, and all $r\in[0,T]$. This implies that $\left|Z_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$ and all $r\in[0,T]$. In particular, we get $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$. Since we chose $L>0$ and $\omega\in\Omega$ arbitrary, $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$. Now, Lemma 4.2 yields that $\lim_{\varepsilon\to 0}R_{1,1}(t,\varepsilon)=0$ and $\lim_{\varepsilon\to 0}R_{1,2}(t,\varepsilon)=0$ in $ucp$.
    Let us next show that $\lim_{\varepsilon\to 0}I_{2}(t,\varepsilon)=\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}\left[W,N\right]_{r}$. We use (2.1) combined with Lemma 4.6 (whose conditions are fulfilled due to the localization arguments at the beginning) in $I_{2}$ to write:
    $I_{2}(t,\varepsilon)$ $=\int_{0}^{t}\int_{r}^{r+\varepsilon}\beta(s,X_{r})\mathrm{d}s\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r+\int_{0}^{t}\int_{r}^{r+\varepsilon}\gamma(s,X_{r})\mathrm{d}W_{s}\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r$
    $=\int_{0}^{t}\int_{r}^{r+\varepsilon}\gamma(r,X_{r})\mathrm{d}W_{s}\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r+\int_{0}^{t}\int_{r}^{r+\varepsilon}\beta(s,X_{r})\mathrm{d}s\displaystyle\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r$
    $\hskip 12.0pt+\int_{0}^{t}\int_{r}^{r+\varepsilon}\left[\gamma(s,X_{r})-\gamma(r,X_{r})\right]\mathrm{d}W_{s}\frac{N_{r+\varepsilon}-N_{r}}{\varepsilon}\mathrm{d}r$
    $=:I_{2,1}(t,\varepsilon)+R_{2,1}(t,\varepsilon)+R_{2,2}(t,\varepsilon).$
    For $I_{2,1}$, we get due to Lemma 4.1 and Assumption 3:
    $I_{2,1}(t,\varepsilon)=\int_{0}^{t}\gamma(r,X_{r})\displaystyle\frac{(W_{r+\varepsilon}-W_{r})(N_{r+\varepsilon}-N_{r})}{\varepsilon}\mathrm{d}r\xrightarrow[ucp]{\varepsilon\to 0}\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}\left[W,N\right]_{r}.$
    For the two terms remaining, we define for $r\in[0,T]$ and $\varepsilon>0$:
    $\widetilde{Y}^{\varepsilon}_{r}$ $:=\displaystyle\frac{1}{\varepsilon}\int_{r}^{r+\varepsilon}\beta(s,X_{r})\mathrm{d}s,$ $\widetilde{Z}^{\varepsilon}_{r}$ $:=\displaystyle\frac{1}{\sqrt{\varepsilon}}\int_{r}^{r+\varepsilon}\left[\gamma(s,X_{r})-\gamma(r,X_{r})\right]\mathrm{d}W_{s}.$
    First, we note that both $\widetilde{Y}^{\varepsilon}$ and $\widetilde{Z}^{\varepsilon}$ are a.s. bounded for all $\varepsilon>0$ since $\sup_{x\in K}|\beta(t,x)|\in\mathcal{L}^{1}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$, $\sup_{x\in K}|\gamma(t,x)|\in\mathcal{L}^{2}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$, and $X$ is a.s. continuous.
    We continue with $R_{2,1}$: In the case of Assumption 1a, classical localization arguments give us the existence of a $\tilde{B}_{1}>0$ such that $\int_{0}^{T}\sup_{s\in[0,T]}|\beta(s,X_{t})|\mathrm{d}t\leq\tilde{B}_{1}$ a.s. and we get a.s.:
    $\int_{0}^{T}|\widetilde{Y}^{\varepsilon}_{r}|\mathrm{d}r\leq\int_{0}^{T}\displaystyle\frac{1}{\varepsilon}\int_{r}^{r+\varepsilon}\left|\beta(s,X_{r})\right|\mathrm{d}s\mathrm{d}r\leq\int_{0}^{T}\sup_{s\in[0,T]}|\beta(s,X_{r})|\mathrm{d}r\leq\tilde{B}_{1}.$
    In the case of Assumption 1b, we get:
    $\mathbb{E}\left[\left|\int_{0}^{T}|\widetilde{Y}^{\varepsilon}_{r}|\mathrm{d}r\right|\right]$ $\leq\mathbb{E}\left[\int_{0}^{T}\displaystyle\frac{1}{\varepsilon}\int_{r}^{r+\varepsilon}\left|\beta(s,X_{r})\right|\mathrm{d}s\mathrm{d}r\right]$
    $=\int_{0}^{T}\displaystyle\frac{1}{\varepsilon}\int_{r}^{r+\varepsilon}\mathbb{E}\left[\left|\beta(s,X_{r})\right|\right]\mathrm{d}s\mathrm{d}r\leq\int_{0}^{T}\sup_{s\in[0,T]}\mathbb{E}\left[\left|\beta(s,X_{r})\right|\right]\mathrm{d}r\leq B_{1},$
    i.e., $\int_{0}^{T}|\widetilde{Y}^{\varepsilon}_{r}|\mathrm{d}r$ is bounded in $\mathcal{L}^{1}(\mathrm{d}\mathbb{P})$. Hence, in both cases it holds that $\int_{0}^{T}|\widetilde{Y}^{\varepsilon}_{r}|\mathrm{d}r$ is bounded in probability and Lemma 4.4 implies that $\lim_{\varepsilon\to 0}R_{2,1}=0$ in $ucp$.
    Finally, we analyze $R_{2,2}$: First, we get for $r\in[0,T]$ fixed:
    $\widetilde{Z}^{\varepsilon}_{r}$ $=\displaystyle\frac{W_{r+\varepsilon}-W_{r}}{\sqrt{\varepsilon}}\cdot\displaystyle\frac{1}{W_{r+\varepsilon}-W_{r}}\int_{r}^{r+\varepsilon}\left[\gamma(s,X_{r})-\gamma(r,X_{r})\right]\mathrm{d}W_{s}.$
[…]
    In the next step, we want to show that this convergence also holds in $L^{2}(\mathrm{d}\mathbb{P})$. For this purpose, we show that $|\widetilde{Z}^{\varepsilon}_{r}|^{2}$ is uniformly integrable. Let $\varrho>0$ be arbitrary. Then, we get using Mao[18, p.39] (with $p=2+\varrho$):
    $\mathbb{E}[|\widetilde{Z}^{\varepsilon}_{r}|^{2+\varrho}]$ $=\displaystyle\frac{1}{\varepsilon^{1+\frac{\varrho}{2}}}\mathbb{E}[\left|\int_{r}^{r+\varepsilon}\left[\gamma(s,X_{r})-\gamma(r,X_{r})\right]\mathrm{d}W_{s}\right|^{2+\varrho}]$
    $\leq\displaystyle\frac{1}{\varepsilon^{1+\frac{\varrho}{2}}}\left(\displaystyle\frac{(2+\varrho)(1+\varrho)}{2}\right)^{\frac{2+\varrho}{2}}\varepsilon^{\frac{\varrho}{2}}\mathbb{E}\left[\int_{r}^{r+\varepsilon}\left|\gamma(s,X_{r})-\gamma(r,X_{r})\right|^{2+\varrho}\mathrm{d}s\right].$
    We define $\tilde{C}:=\left(\frac{(2+\varrho)(1+\varrho)}{2}\right)^{\frac{2+\varrho}{2}}$ and $C:=\tilde{C}\cdot 2^{2+\varrho}$. In particular, $\tilde{C}\leq C$. Now, we get for all $r\in[0,T]$ using Lemma 4.7:
    $\mathbb{E}[|\widetilde{Z}^{\varepsilon}_{r}|^{2+\varrho}]$ $\leq\tilde{C}\cdot\displaystyle\frac{1}{\varepsilon}\int_{r}^{r+\varepsilon}\mathbb{E}\left|\gamma(s,X_{r})-\gamma(r,X_{r})\right|^{2+\varrho}\mathrm{d}s$
    $\leq\tilde{C}\sup_{s\in[0,T]}\mathbb{E}[\left|\gamma(s,X_{r})-\gamma(r,X_{r})\right|^{2+\varrho}]\leq C\sup_{s\in[0,T]}\mathbb{E}[\left|\gamma(s,X_{r})\right|^{2+\varrho}]\leq C\cdot B_{2}.$ (4.1)
    Thus, $\widetilde{Z}^{\varepsilon}_{r}\xrightarrow[\mathcal{L}^{2}(\mathrm{d}\mathbb{P})]{\varepsilon\to 0}0$ for all $r\in[0,T]$. Analogously to (Proof of Theorem 3.1.), we also get $\sup_{r\in[0,T]}\mathbb{E}[|\widetilde{Z}^{\varepsilon}_{r}|^{2}]\leq C\cdot B_{2}$. Hence, the bounded convergence theorem gives us that $\int_{0}^{T}\mathbb{E}|\widetilde{Z}^{\varepsilon}_{r}|^{2}\mathrm{d}r\xrightarrow{\varepsilon\to 0}0$. Thus, $\int_{0}^{T}|\widetilde{Z}^{\varepsilon}_{r}|^{2}\mathrm{d}r\xrightarrow[\mathbb{P}]{\varepsilon\to 0}0$ and Lemma 4.5 implies that $\lim_{\varepsilon\to 0}R_{2,2}=0$ in $ucp$ which gives us the proof of the Theorem. ∎
5 Applications
In this section, we apply Theorem 3.1 to derive a novel stochastic calculus formula for quadratic variations and a representation result for strong solutions of SPDEs.
5.1 Formulas for quadratic variations
In this subsection, we give in Proposition 5.1 a novel stochastic calculus formula for quadratic variations. Afterwards, we point out some special cases. Note that we cannot use the Itô-Ventzell formula in $C^{0,2}$ or $C^{1,2}$ since the function $F$ defined below is not twice continuously differentiable in the space direction. We remark further that Proposition 5.1 is not a special case of the Meyer-Itô formula (see, for instance, Protter [19]) which is only applicable for delta-convex functions, i.e., for the difference of two continuous convex functions. However, the class of $C^{1}$-functions and the class of delta-convex functions are not included in each other, see, e.g., Zelený [27].
Proposition 5.1.
Let $W$ be a one-dimensional Brownian Motion, $X=M+A$ be a continuous weak Dirichlet FQV process, and $f:\Omega\times[0,T]\times\mathbb{R}\to\mathbb{R}$ be such that $f(\cdot,x)$ is previsible for all $x\in\mathbb{R}^{d}$ fixed, and continuously differentiable in $x$ for all $t\in[0,T]$. Moreover, we assume that for all $K\subset\mathbb{R}$ compact exists a $C_{K}(\omega)\geq 0$ with $\mathbb{E}[C_{K}]<\infty$ such that $\sup_{(t,x)\in[0,T]\times K}(|f(\omega,t,x)|+|f_{x}(\omega,t,x)|)\leq C_{K}(\omega)$. Then, it holds that
$\left[\int_{0}^{t}f(r,x)\mathrm{d}W_{r}\big{|}_{x=X_{t}},W_{t}\right]=\int_{0}^{t}f(r,X_{r})\mathrm{d}r+\int_{0}^{t}\int_{0}^{r}f_{x}(s,x)\mathrm{d}W_{s}\big{|}_{x=X_{r}}\mathrm{d}[M,W]_{r}.$
Proof. First, we define $F(t,x):=\int_{0}^{t}f(r,x)\mathrm{d}W_{r}$ (where we suppress the $\omega$ as usual). Then, $F$ is a stochastic flow in the sense of Definition 2.7 with local characteristics $\beta(t,x)=0$ and $\gamma(t,x)=f(t,x)$. The assumptions of Theorem 3.1 are fulfilled since continuous differentiability implies local Lipschitz continuity. Moreover, it holds that $F_{x}(t,x)=\int_{0}^{t}f_{x}(r,x)\mathrm{d}W_{r}$ due to the stochastic Leibniz integral rule. Then, Theorem 3.1 implies with the decomposition $X=M+A$:
$\int_{0}^{t}f(r,x)\mathrm{d}W_{r}\big{|}_{x=X_{t}}=0+\int_{0}^{t}f(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\int_{0}^{r}f_{x}(s,x)\mathrm{d}W_{s}\big{|}_{x=X_{r}}\mathrm{d}M_{r}+\mathcal{B}^{X}(F)_{t},$
where $\mathcal{B}^{X}(F)_{t}$ is a weak zero energy process. Hence, we get for any continuous local martingale $\tilde{M}$:
$\left[\int_{0}^{t}f(r,x)\mathrm{d}W_{r}\big{|}_{x=X_{t}},\tilde{M}_{t}\right]=\int_{0}^{t}f(r,X_{r})\mathrm{d}[W,\tilde{M}]_{r}+\int_{0}^{t}\int_{0}^{r}f_{x}(s,x)\mathrm{d}W_{s}\big{|}_{x=X_{r}}\mathrm{d}[M,\tilde{M}]_{r}.$
Finally, choosing $\tilde{M}:=W$ yields the proposition since $\mathrm{d}[W]_{r}=\mathrm{d}r$.
Corollary 5.2.
Let $W$ be a one-dimensional Brownian Motion, and $f:\mathbb{R}\to\mathbb{R}$ be a continuously differentiable function. Then, we get that:
$[f(W_{t})W_{t},W_{t}]=\int_{0}^{t}f(W_{r})\mathrm{d}r+\int_{0}^{t}f^{\prime}(W_{r})W_{r}\mathrm{d}r.$
Proof. This corollary follows immediately from Proposition 5.1 after setting $X=W$ with $\mathrm{d}[W]_{r}=\mathrm{d}r$. Moreover, note that $\int_{0}^{t}\int_{0}^{r}f_{x}(x)\mathrm{d}W_{s}\big{|}_{x=X_{r}}\mathrm{d}[M,W]_{r}=\int_{0}^{t}f^{\prime}(X_{r})W_{r}\mathrm{d}[M,W]_{r}$ in this setting.
For Corollary 5.2, we could have alternatively tried to use an Itô-formula with weaker differentiability conditions.
5.2 A representation result for strong solutions of SPDEs
In this subsection, the overall target is to give a representation result for strong solutions of stochastic elliptic PDE for a special type of weak Dirichlet processes. We start by giving a generalization of the solution theory of SPDEs (naturally generalizing the theory of Gozzi and Russo [13] who considered PDEs).
Let $b:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ be two continuous functions. Then, we define the time-dependent elliptic operator $L_{0}(t):D(L_{0})\subset C^{0}([0,T]\times\mathbb{R}^{d})\to C^{0}([0,T]\times\mathbb{R}^{d})$ with $D(L_{0})=C^{0,2}([0,T]\times\mathbb{R}^{d})$ by $L_{0}(t)F(t,x):=b(t,x)F_{x}(t,x)+\frac{1}{2}tr(\sigma^{\intercal}(t,x)F_{xx}(t,x)\sigma(t,x)$, where $tr$ denotes the trace of a matrix. We then consider the stochastic inhomogeneous time-dependent elliptic problem for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$ and almost all $\omega\in\Omega$:
$L_{0}(t)F(\omega,t,x)=h(\omega,t,x),$ (5.1)
where $h(\omega,\cdot)\in C^{0}([0,T]\times\mathbb{R}^{d})$ almost surely.
Definition 5.3.
We say that $F$ with local characteristics $(\beta(t,x),\gamma(t,x))$ is a strict solution to the stochastic elliptic problem (5.1) if for a.e. $\omega\in\Omega$ it holds that $F(\omega,\cdot)\in D(L_{0})$, $\beta(\omega,\cdot)\in C^{0,1}([0,T]\times\mathbb{R})$, $\gamma(\omega,\cdot)\in C^{0,1+\alpha}([0,T]\times\mathbb{R}^{d})$ for an $\alpha\in(0,1)$, and (5.1) holds.
Analogously to Gozzi and Russo [13] who defined a strong solution for PDEs, we define a strong solution for SPDEs as follows:
Definition 5.4.
We say that $F$ is a strong solution to the stochastic elliptic problem (5.1) if there exists a sequence $(F^{n})_{n\in\mathbb{N}}$ with local characteristics $(\beta^{n}(\omega,t,x),\gamma^{n}(\omega,t,x))$ and a sequence $(h^{n})_{n\in\mathbb{N}}$ such that for every $n\in\mathbb{N}$ and a.e. $\omega\in\Omega$ holds:
  1. 1
    $F^{n}(\omega,\cdot)\in D(L_{0})$, and $h^{n}(\omega,\cdot)\in C^{0}([0,T]\times\mathbb{R}^{d})$,
  2. 2
    $\beta^{n}(\omega,\cdot)\in C^{0,1}([0,T]\times\mathbb{R})$, and $\gamma^{n}(\omega,\cdot)\in C^{0,1+\alpha}([0,T]\times\mathbb{R}^{d})$ for an $\alpha\in(0,1)$,
  3. 3
    $F^{n}$ is a strict solution of $L_{0}(t)F^{n}(\omega,t,x)=h^{n}(\omega,t,x)$ for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$ and a.e. $\omega\in\Omega$,
  4. 4
    $F^{n}(\omega,\cdot)\xrightarrow{n\to\infty}F(\omega,\cdot)$ and $h^{n}(\omega,\cdot)\xrightarrow{n\to\infty}h(\omega,\cdot)$ in $C^{0}([0,T]\times\mathbb{R}^{d})$ (equipped with the supremum-norm on compact sets), and
  5. 5
    $\beta^{n}(\omega,\cdot)\xrightarrow{n\to\infty}\beta(\omega,\cdot)$ in $\mathcal{L}^{1}(K;\mathrm{d}t)$ for all compact sets $K\subset[0,T]\times\mathbb{R}$.
Note that this a generalization of the notion of a strict solution since, of course, every strict solution is also a strong solution by taking the constant sequences $F^{n}:=F$ and $h^{n}:=h$. Our representation theorem states then:
Theorem 5.5.
Let $b:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ be two continuous functions, and $F$ (with local characteristics $(\beta(\omega,t,x),\gamma(\omega,t,x))$) be a strong solution of problem (5.1) where we denote the approximation sequence by $F^{n}$ (with local characteristics $(\beta^{n}(\omega,t,x),\gamma^{n}(\omega,t,x))$). Define $X_{t}:=x_{0}+\int_{0}^{t}\sigma(r,X_{r})\mathrm{d}W_{r}+A_{t}$, where $x_{0}\in\mathbb{R}^{d}$ and $A$ is a continuous weak zero energy process such that $X$ is a continuous FQV process. Furthermore, we assume:
  1. 1
    $F(\omega,\cdot)\in C^{0,1}([0,T]\times\mathbb{R}^{d})$ for a.e. $\omega\in\Omega$,
  2. 2
    $\beta$ and $\gamma$ fulfill the assumptions of Theorem 3.1 with $\gamma(\omega,\cdot)\in C^{0,1}([0,T]\times\mathbb{R}^{d})$,
  3. 3
    It holds in $ucp$ that
    $\lim_{n\to\infty}\int_{0}^{t}(F_{x}(r,X_{r})-F^{n}_{x}(r,X_{r}))\mathrm{d}^{-}A_{r}$ $+\int_{0}^{t}\mathrm{tr}((\gamma_{x}(r,X_{r})-\gamma^{n}_{x}(r,X_{r}))\sigma(r,X_{r}))\mathrm{d}r$
    $+\int_{0}^{t}(F_{x}(r,X_{r})-F^{n}_{x}(r,X_{r}))b(r,X_{r})\mathrm{d}r=0.$ (5.2)
Then, we have that (suppressing the dependence on $\omega$)
$F(t,X_{t})=$ $F(0,X_{0})+\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F_{x}(r,X_{r})\sigma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta(r,X_{r})\mathrm{d}r$
$+\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\mathrm{tr}(\gamma_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r+\int_{0}^{t}h(r,X_{r})\mathrm{d}r$
$-\int_{0}^{t}b(r,X_{r})F_{x}(r,X_{r})\mathrm{d}r.$
Proof. Since $F$ is a strong solution of (5.1) with approximating sequences $F^{n}$ and $h^{n}$, we get with the second part of Theorem 3.1 (suppressing the dependence on $\omega$):
$F^{n}(t,X_{t})=$ $F^{n}(0,X_{0})+\int_{0}^{t}\gamma^{n}(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F^{n}_{x}(r,X_{r})\sigma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta^{n}(r,X_{r})\mathrm{d}r$
$+\int_{0}^{t}F^{n}_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\mathrm{tr}(\gamma^{n}_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r$
$+\frac{1}{2}\int_{0}^{t}\mathrm{tr}(\sigma^{\intercal}(r,X_{r})F^{n}_{xx}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r$
$=$ $F^{n}(0,X_{0})+\int_{0}^{t}\gamma^{n}(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F^{n}_{x}(r,X_{r})\sigma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta^{n}(r,X_{r})\mathrm{d}r$
$+\int_{0}^{t}F^{n}_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\mathrm{tr}(\gamma^{n}_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r+\int_{0}^{t}L_{0}(r)F^{n}(r,X_{r})\mathrm{d}r$
$-\int_{0}^{t}b(r,X_{r})F^{n}_{x}(r,X_{r})\mathrm{d}r.$ (5.3)
Now, we define:
$M_{t}^{n}:=$ $F^{n}(t,X_{t})-F^{n}(0,X_{0})-\int_{0}^{t}\beta^{n}(r,X_{r})\mathrm{d}r-\int_{0}^{t}F^{n}_{x}(r,X_{r})\mathrm{d}^{-}A_{r}-\int_{0}^{t}\mathrm{tr}(\gamma^{n}_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r$
$-\int_{0}^{t}L_{0}(r)F^{n}(r,X_{r})\mathrm{d}r+\int_{0}^{t}b(r,X_{r})F^{n}_{x}(r,X_{r})\mathrm{d}r,$
$M_{t}:=$ $F(t,X_{t})-F(0,X_{0})-\int_{0}^{t}\beta(r,X_{r})\mathrm{d}r-\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}-\int_{0}^{t}\mathrm{tr}(\gamma_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r$
$-\int_{0}^{t}h(r,X_{r})\mathrm{d}r+\int_{0}^{t}b(r,X_{r})F_{x}(r,X_{r})\mathrm{d}r.$ (5.4)
Note that for almost all $\omega\in\Omega$, it holds that $L_{0}(r)F^{n}(r,X_{r}(\omega))=h^{n}(r,X_{r}(\omega))\xrightarrow{n\to\infty}h(r,X_{r}(\omega))$ uniformly in $r$ due to 4 in Definition 5.4 and $(r,X_{r}(\omega))$ taking values in a compact set for $r\in[0,t]$. Then, by the assumptions of a strong solution and by (3), it holds that $M_{t}^{n}\xrightarrow{n\to\infty}M_{t}$ in $ucp$. Moreover, it follows from (5.3) that $M_{t}^{n}$ is a local martingale. Proposition 4.4 of Gozzi and Russo [13] implies then that $M$ is also a local martingale. Furthermore, note that $M_{t}$ is continuous in $t$ since $X$, $F$, and $A$ are continuous. Then, we define the weak zero energy process
~$\tilde{\mathcal{B}}^{X}(F)_{t}:=$ $\int_{0}^{t}\beta(r,X_{r})\mathrm{d}r+\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\mathrm{tr}(\gamma_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r+\int_{0}^{t}h(r,X_{r})\mathrm{d}r$
$-\int_{0}^{t}b(r,X_{r})F_{x}(r,X_{r})\mathrm{d}r.$
Rearranging (5.4) implies that $F(t,X_{t})=F(0,X_{0})+M_{t}+\tilde{\mathcal{B}}^{X}(F)_{t}$. Next, we apply the first part of Theorem 3.1 on $F$ to get that $F(t,X_{t})=F(0,X_{0})+\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F_{x}(r,X_{r})\sigma(r,X_{r})\mathrm{d}W_{r}+\mathcal{B}^{X}(F)_{t}$, where $\tilde{M}_{t}:=\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F_{x}(r,X_{r})\sigma(r,X_{r})\mathrm{d}W_{r}$ is a continuous local martingale. Finally, since $M_{t}$ and $\tilde{M}_{t}$ are continuous, the decomposition of the weak Dirichlet process $F(t,X_{t})$ is unique. In particular, we get that $M_{t}=\tilde{M}_{t}$ and $\mathcal{B}^{X}(F)_{t}=\tilde{\mathcal{B}}^{X}(F)_{t}$ which implies the claim.
5.3 An example from financial mathematics
In this subsection, we consider a simplified version of a model introduced in Bank and Baum [3] which describes a wealth process of a large investor with reduced differentiability assumptions. The large investor directly influences the price with the amount of shares she/he buys/sells.
For this purpose, let $T>0$ and define for $0\leq t\leq T$ a price fluctuation process $P_{t}^{\vartheta}=P(t,\vartheta)$, $\vartheta\in\mathbb{R}$, of the risky asset as a stochastic flow with local characteristics $(\beta(t,x),\gamma(t,x))$ as in Definition 2.7 with $\beta(t,x)\in C^{0,0}$ and $\gamma(t,x)\in C^{0,\alpha}$ for an $\alpha\in(0,1)$. The parameter $\vartheta$ denotes the number of shares the large investor holds. Moreover, we assume that $P$ fulfills the assumptions of Theorem 3.1. (This is a relaxation compared to Bank and Baum [3], where second order differentiability is needed.) Let $X=(X_{t})_{t\in[0,T]}$ be a (self-financingThat is, assuming that there are no intermediate capital injections.) strategy for the large investor describing the number of shares held. Assuming that $X$ is a semimartingale, according to Bank and Baum [3, p.5], the discounted bank account value, say $\Xi^{X}$, of the large investor evolves as
$\Xi_{t}^{X}=\Xi_{0}-\int_{0}^{t}P(r,X_{r})\mathrm{d}X_{r}-[P(\cdot,X),X]_{t},\quad\text{for $0\leq t\leq T$,}$ (5.5)
where the integral models the fluctuations of the portfolio value until time $t$ and the quadratic variation term accounts for the sensitivity of the prices regarding the large investor’s orders at time $t$. By splitting an order into smaller packages, Bank and Baum [3] argue that when liquidating the portfolio, the amount obtained by the large investor is given by $L(t,\vartheta)=\int_{0}^{\vartheta}P(t,x)\mathrm{d}x$, where $\vartheta$ denotes the large investor’s number of shares before liquidation. The real (or asymptotically realizable) wealth $V^{X}$ which can be achieved by the strategy $X$ until time $t$ is consequently given by
$V_{t}^{X}=\Xi_{t}^{X}+L(t,X_{t}).$ (5.6)
Proposition 5.6.
Let $X$ be a continuous self-financing strategy which is a semimartingale. The dynamics of the real wealth process $V_{t}^{X}$ is then given by:
$V_{t}^{X}=V_{0}^{X}+\int_{0}^{t}\beta^{L}(r,X_{r})\mathrm{d}r+\int_{0}^{t}\gamma^{L}(r,X_{r})\mathrm{d}X_{r}-\frac{1}{2}\int_{0}^{t}P_{x}(r,X_{r})\mathrm{d}[X]_{r},$
where $\beta^{L}(t,\vartheta)=\int_{0}^{\vartheta}\beta(t,x)\mathrm{d}x$ and $\gamma^{L}(t,\vartheta)=\int_{0}^{\vartheta}\gamma(t,x)\mathrm{d}x$.
Proof. First, we can use the standard decomposition $X=M+A$ with $M$ being a martingale and $A$ a finite variation process since $X$ is a semimartingale. Next, we notice by the definition of $L$, that $L$ is also a stochastic flow with local characteristics $(\beta^{L}(t,x),\gamma^{L}(t,x))$ in the sense of Definition 2.7 and fulfills the assumption of Theorem 3.1 (b). Indeed, we are allowed to interchange the integrals when determining the local characteristics of $L$ since $\beta(s,x)\in C^{0,0}$ and thus $\beta$ is bounded on $[0,t]\times[0,\vartheta]$ for a.e. $\omega\in\Omega$ fixed. Furthermore, $\gamma(s,x)\in C^{0,\alpha}$, and we can therefore apply the stochastic Fubini theorem, given in Lemma 2.6 resp. Lemma 2.7 in Krylov [16]), for the stochastic integral. Hence, we get with Theorem 3.1 (b):
$L(t,X_{t})=$ $L(0,X_{0})+\int_{0}^{t}\beta^{L}(r,X_{r})\mathrm{d}r+\int_{0}^{t}\gamma^{L}(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}L_{x}(r,X_{r})\mathrm{d}M_{r}$
$+\int_{0}^{t}L_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\gamma^{L}_{x}(r,X_{r})\mathrm{d}[X,W]_{r}+\frac{1}{2}\int_{0}^{t}L_{xx}(r,X_{r})\mathrm{d}[X]_{r}$
$=$ $L(0,X_{0})+\int_{0}^{t}\beta^{L}(r,X_{r})\mathrm{d}r+\int_{0}^{t}\gamma^{L}(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}P(r,X_{r})\mathrm{d}X_{r}$
$+\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}[X,W]_{r}+\frac{1}{2}\int_{0}^{t}P_{x}(r,X_{r})\mathrm{d}[X]_{r},$ (5.7)
since $L_{x}=P$ with $\gamma_{x}^{L}=\gamma$. Next, we apply Theorem 3.1 (a) to $P$ which gives us:
$P(t,X_{t})=$ $P(0,X_{0})+\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}P_{x}(r,X_{r})\mathrm{d}M_{r}+\mathcal{B}^{X}(P)_{t}.$
Hence, it follows that:
$[P(\cdot,X),X]_{t}$ $=\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}[X,W]_{r}+\int_{0}^{t}P_{x}(r,X_{r})\mathrm{d}[X,M]_{r}$
$=\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}[X,W]_{r}+\int_{0}^{t}P_{x}(r,X_{r})\mathrm{d}[X]_{r},$
since $[A,M]_{r}=0=[A]_{r}$ due to $A$ being a finite variation process. Thus, we get for the bank account value $\Xi$ (defined in (5.5)):
$\Xi_{t}^{X}=\Xi_{0}-\int_{0}^{t}P(r,X_{r})\mathrm{d}X_{r}-\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}[X,W]_{r}-\int_{0}^{t}P_{x}(r,X_{r})\mathrm{d}[X]_{r}.$ (5.8)
Next, we plug (5.7) and (5.8) into (5.6), then several terms cancel out and the claim follows.